### v1

# Constant search time algorithm via topological quantum walks

Abstract
It is well-known that quantum algorithms such as Grover’s can provide a quadradic speed-up for unstructured search problems. By adding topological structure to a search problem, we show that it is possible to achieve a constant search-time quantum algorithm with a constant improvement of the search probability over classical search. Specifically, we study the spatial search algorithm implemented by a two-dimensional split-step quantum random walks that realize topologically nontrivial phases and show the asymptotic search behavior is constant with growing system size. Using analytical and numerical calculations, we determine the efficient search regions in the parameter space of the quantum walker. These regions correspond to pairs of trapped states formed near a lattice defect. By studying the spectral properties of the discrete time-evolution-operators, we show that these trapped states have large overlap with the initial state. This correspondence, which is analogous to localization by constructive interference of bound states, makes it possible to reach the best possible search-time asymptotic and produce a disorder-protected fast search in quantum random walks.
Quantum search algorithm, Grover’s search [1], is known as an algorithm that demonstrates quantum advantage over its classical analogue. Grover’s algorithm performs a search of marked element in an unstructured database with a given probability with asymptotic time being a square root of database element count. However, its implementation depends on an oracle operator - a way of distinguishing marking element that applies to a whole quantum state. The quantum random walks are known as an example of natural implementation of such an oracle operator [2, 3, 4, 5, 6]. The discrete-time quantum random walks are defined as a set of shift and spin flip operators applied at the whole system at each time step. The successful search event is defined as a time of maximum localization of quantum walker on a defect when starting from a uniform distribution over all system.
Although, different kinds of quantum walks demonstrate a Grover-type convergence of a spatial search in two and higher dimensions, the dependence on spin flip operator and position of defect site relative to remaining system is crucial and in many cases results in a decreased efficiency of search algorithm [5, 4, 7]. The latter fact also serves as a motivation for the search for sub-optimal yet more robust to disorder and element positioning algorithms, that would provide quantum advantage. In the present paper we study the two-dimensional split-step quantum random walks that are analogues of Chern topological insulators [8] in crystalline solids. The similar split-step QRW in 1D were studied in Refs.[9, 10] and a number of our conclusions made below might work also in those cases. By analyzing the parameter spaces of quantum walker and defect, we find that there exist sets of parameters that demonstrate efficient search of a given defect (see a search probability map in Fig.1). Below we show that these regions in parameter space can be categorized into two distinct classes based on their position inside the topological phase or on the phase separation lines. Notably, the search time for the parameters inside both classes of regions demonstrates saturation to a constant search time with growing system size. The saturation of search time happens at different system sizes, while before that the square root time growth behavior takes place.
Figure 1: The success probability of search algorithm as a function of QRW parameters $\Theta_{1,2}$ (see Eq.(2)) for a defect with $\Theta_{1}^{def}=5\pi/8$, $\Theta_{2}^{def}=\pi/2$. High peaks are of order of magnitude larger than average localization probability and correspond to parameters for which quantum walker performs search efficiently. Inset: geometry of split-step QRW on a triangular lattice with three shift vectors ${\boldsymbol{v}}_{i}$ solid ones correspond to shift of spin up components, dashed $-{\boldsymbol{v}}_{i}$ for spin down.
The constant search time asymptotic was previously indentified only for electric Dirac QRW with non-local marking of defect in the form of Coulomb potential [11, 12]. The non-locality creates an important difference from our algorithm. Namely, here the constant search time is achieved with local marking, thus making it directly implementable in experimental memory setup.
Below we analyze in detail the time evolution of states for a range of quantum walker parameters and study the overlap of initial state with the eigenstates of a unitary operator corresponding to a single time step of QRW. We demonstrate that the search speedup with constant time asymptotic is related to the appearance of trapped states of a specific structure near the defect: the two pairs of trapped states should simultaneously have a large matrix element of coordinate operator of the defect and the large overlap with initial uniformly distributed state. In the mathematical sense the trapped states are defined as a states that have compact support and zero density elsewhere. Their analogue in quantum mechanics are the strongly localized bound states on impurity in a system. We note that a number of mathematically rigorous studies have been made for spectral properties of quantum random walks and trapped states [5, 13, 9, 14, 15, 16] together with their dynamics [17, 18]. However, the possible efficiency of search algorithm based on exploiting the particular structure of QRW was not previously addressed. The strong dependence of the trapped states structure on defect and walker parameters also suggests a scenario of protecting quantum search from a random disorder by an appropriate fine-tuning of marking parameters for the searched element.
The quantum random walk deployed for algorithmic speed-up in this work was firstly introduced in Ref.[8] as a way to simulate topological insulators. Its single time step
$\Psi(t+1)=U(\Theta_{1},\Theta_{2})\Psi(t)$ (1)
is defined via three rotations of spin and subsequent translations:
$U(\Theta_{1},\Theta_{2})=T_{{\boldsymbol{v}}_{3}}R\left(\theta_{1}\right)T_{{\boldsymbol{v}}_{2}}R\left(\theta_{2}\right)T_{{\boldsymbol{v}}_{3}}R\left(\theta_{1}\right),$ (2)
$R(\theta)=\left[\begin{array}[]{cc}\cos(\theta/2)&-\sin(\theta/2)\\ \sin(\theta/2)&\cos(\theta/2)\end{array}\right].$ (5)
The translation operations shift spin up components by $+{\boldsymbol{v}}_{i}$ and spin down by $-{\boldsymbol{v}}_{i}$. The vectors ${\boldsymbol{v}}_{1,2,3}$ can be defined on a triangular lattice as $v_{1}=(1,0)$, $v_{2}=(1/2,\sqrt{3/2})$ and $v_{3}=(1/2,-\sqrt{3/2})$, see inset in Fig.1.
[…]
Figure 2: (a) A colormap representation of a maximum localization probability on a defect site over the evolution time $T=1000$ for a system of size $N=40\times 40$ starting from a fully uniform distribution. Black numbers indicate Chern numbers of the corresponding parameter regions and dashed lines correspond to phase separation lines. Cyan dot marks the defect parameters $(5\pi/8,\pi/2)$. (b) Search time measured as a position of first probability peak as a function of system size for walker parameters indicated by crosses ‘1’ and ‘2’ in panel (a). (c) Search probability dependence on system size. The two regions of different dependence are present where probability follows Grover-type scaling $P\sim 1/\log(N)$ until the time asymptotic changes to constant in (b) and then switches to system size scaling $P\sim 1/N$ but having few orders of magnitude larger value than average density in the system. (d) Comparison of performance with Grover’s algorithm and repeated classical search for system size $\sqrt{N}=512$. Walker $2$ shows advantage over classical.
In our setup the search algorithm goal is defined as a problem of finding a single node on a finite 2D lattice of size $\sqrt{N}\times\sqrt{N}$ unit cells starting from initial state $|i\rangle$. We apply the periodic boundary condition in order to avoid the appearance of edge states. The searched node is marked by parameters $\Theta_{1,2}^{def}$ that are different from surrounding $\Theta_{1,2}$. To estimate the possibility of given quantum walker to perform search efficiently, we firstly scan the entire parameter range $\Theta_{1,2}$ for a given defect values $\Theta_{1,2}^{def}$. The performance of search algorithm is measured by a search probability. It is defined as a peak $\max_{t\in[0,T]}P_{def}(t)$ in time-dependent density of the system state at the defect node,
$P_{def}(t)=|\langle d|\Psi(t)\rangle|^{2},$ (6)
where $|d\rangle=\sum_{\sigma=\uparrow,\downarrow}|x=def,\sigma\rangle$ is the defect coordinate operator with both components of spin $\sigma$, and $|i\rangle$ is the initial state of the system defined as $|i\rangle=\frac{1}{\sqrt{2N}}\sum_{\sigma,x}|x,\sigma\rangle$ with summation over entire system. The plot of the maximum value of density $P_{def}(t)$ on defect over the interval of time $T=1000$ for test system size $N=40\times 40$ sites is shown in Fig.1 for $\Theta_{1,2}^{def}=(5\pi/8,\pi/2)$. In Fig.2(a) we show the colormap representation of the same search probability distribution on the top of phase diagram in parameter space (see also Supplement [21] for other defect parameters).
The islands of efficient search in parameter space fill a small areas placed in two distinct types of regions: inside topological phases and on phase separation lines. We compare the search time and probability behavior with growing system size for the walker parameters marked by ‘1’ and ‘2’ in the Fig.2(a), that realize best search probability within its regions. Such results are shown in Fig.2(b). With the growing system size the behavior of search time defined via first probability peak demonstrates qualitative change - from $T\sim\sqrt{N}$ dependence to $T\sim const$ in both curves. As we noted previously, such speed-up was observed only for the Dirac QRW with particular type of Coulomb-like disorder potential [11, 12]. In Fig.2(c) the search probability dependence of walkers ‘1’ and ‘2’ on system size is compared with average density at non-defect nodes and Grover’s algorithm search probability. We find that the constant search time regime corresponds to the $P_{def}(t)=O(1/N)$ with a large numerical prefactor. This decrease of search probability with system size $N$ is faster than Grover’s one, $P_{def}^{Grover}=O(1/log\,N)$. The search probability dependence in the form $O(1/N)$ arises due to the trapped states having finite support. The system size grows beyond the trapped state support size corresponds to saturation of search time asymptotic to constant and clear change in probability dependence curve in Fig.2(c). To check the dependence of search probability on defect parameters for given walker we performed the same 2D parameter space calculation taking walker "$1"$ and checking $\Theta_{1,2}^{def}\in[-2\pi,2\pi]$. It is found that the search probability does not depend on $\Theta_{2}^{def}$ [21], thus the more informative is to present a dependence on $\Theta_{1}^{def}$ as shown in Fig.3(a).
The asymptotic scaling of $O(1/N)$ in search probability does not reach the optimal Grover scaling, but still provides constant improvement over the classical search algorithm. The probability of classical algorithm grows as $t/N$ with multiple trials, $t$ (one per time step). For this purpose, in Fig.2(d) we present a comparison of probability on searched node as a function of time for two walkers marked in Fig.2(a), Grover’s search for split-step QRW from Ref.[7] and a classical search. The walker ‘2’ demonstrates twice as large probability as classical search, while the walker ‘1’ follows the classical search probability for approximately 100 steps. In terms of detection via measurements, higher probability on a node means that the detector resolution used to identify position could be lower - in particular, for walker ‘1‘ the detection resolution could be $100$ times smaller that for classical probes of individual nodes. The Grover’s algorithm performance demonstrates even higher probability at the same time scale, which makes it advantageous in the non-disordered system. As we point out below, that advantage of Grover’s search might be lost in a disordered systems.
To describe the origin of different speed-ups in this system and why the regions of efficient search represent a small fraction of parameter space, we analyze the spectrum and localization properties of eigenvectors of a unitary operator of system with one defect. The probability evolution at a defect site at discrete times $t$ can be rewritten through the unitary operator of QRW,
$P_{def}(t)=|\langle d|\bar{U}^{t}(\Theta_{1},\Theta_{2})|i\rangle|^{2}.$ (7)
The $\bar{U}$ notation is introduced in order to underline that unitary operator takes into account defect site. Substituting the eigenbasis decomposition of the $\bar{U}$ operator, $\bar{U}|n,\pm\rangle=e^{i\pm E_{n}}|n,\pm\rangle$, we have
$P_{def}(t)=\bigg{|}\langle d|\sum_{n,\lambda=\pm}e^{i\lambda E_{n}t}|n,\lambda\rangle\langle n,\lambda|i\rangle\bigg{|}^{2}$ (8)
The saturation of search time to constant scaling after system size becomes large enough suggests that the properties of states that mainly contribute to the Eq.(8) do not change with system size. This feature directly corresponds to the definition of trapped states previously introduced in Refs. [5, 13] in the context of QRWs staying localized around its initial starting node over infinite evolution time. We investigate the contribution of different eigenstates of $\bar{U}$ to the evolution of probability, $P(t)$, in Fig.3(b) by comparing the product of overlaps $|\langle d|n,\lambda\rangle\langle n,\lambda|i\rangle|^{2}$. The panels in Fig.3(b) correspond to different parameters of defect denoted by vertical dashed lines for walker `$1^{\prime}$ in Fig.3(a).
Figure 3: (a) Comparison of maximum value of probability on defect for a walker ‘1’ from Fig.2 as a function of defect parameter $\Theta_{1}^{def}$ with the maximal value of squared of overlaps defined in Eq.(9) with $\mathcal{M}$ containing one and two states respectively. All quantities are normalized to scale (0,1) for the purpose of comparison. The summation over two trapped states gives very accurate approximation of search probability peak. (b) The squared product of overlaps plotted as a function of energy of each eigenstate of operator $\bar{U}$ for walker parameters ‘1’ and defect parameter $\Theta_{1}$ indicated on top of each panel. Trapped states differ from other states by significantly higher value of overlap.
The scale of $x$-axis in each panel of Fig.3(b) shows that defect parameters with higher search probability also correspond to higher overlap product for a few states - in this case two pairs of particle-hole symmetric states. To further verify this observation, we compare the maximal value of overlap product
$W_{\mathcal{M}}=|\sum_{n\in\mathcal{M},\lambda=\partial}\langle d|n,\lambda\rangle\langle n,\lambda|i\rangle|^{2}$ (9)
of a single eigenstate of $\bar{U}$ with the normalized probability dependence $max[P_{def}(t)]$ as function of $\Theta_{1}^{def}$. The comparison is made by taking a single pair of states that maximize overlap with set of indices =$\mathcal{M}=\{n_{max}\}$, as well as two pairs of states =$\mathcal{M}=\{n_{max},n_{second\,max}\}$. The results presented in Fig.3(a) indicate that while =$\mathcal{M}=\{n_{max}\}$ selection overestimates the region of good search, the =$\mathcal{M}=\{n_{max},n_{second\,max}\}$ very precisely describes the high search probability peak. To further check this correspondence between the evolution of described by two pairs of states with maximal overlap and the max probability distribution, we plotted the same comparison for a defect $\Theta_{1,2}^{def}=(5\pi/8,\pi/2)$ along phase separation line and the line with walker `$1^{\prime}$ in Supplemental material. The correspondence for a test system size with $\sqrt{N}=40$ shows that efficient search for such split-step QRW is mostly determined by the appearance of two pairs of trapped states having large overlap with both defect and initial state.
Finally, we perform a comparison of time evolution defined by two pairs of trapped states that maximize overlap criteria with the evolution of probability on the defect. For this purpose we note that the $\bar{U}$ has the same particle-hole symmetry as $U(\Theta_{1,2})$ because it is a real matrix. In addition, by calculating overlap products denoted as $w_{\pm,j}^{\sigma}=\langle d,\sigma|j\rangle\langle j|i\rangle$ for each spin $\sigma=\uparrow,\downarrow$ and pair of trapped states $j$ at the defect site separately, we find that for each trapped state there is a phase difference $w_{+,j}^{\downarrow}=iw_{+,j}^{\uparrow}$ and corresponding $w_{-,j}^{\downarrow}=-iw_{-,j}^{\uparrow}$ up to numerical precision. The former symmetry is present only at defect site and is a property of trapped states localized around the defect. Thus, evolution of both spin components at the defect site that is generated by trapped states is described as follows
$\Psi_{\uparrow,d}(t)=2\sum_{j=1,2}\left|w_{+,j}^{\uparrow}\right|\cos\left(E_{j}t+\arg\left(w_{+,j}^{\uparrow}\right)\right),$ (10)
$\Psi_{\downarrow,d}(t)=2\sum_{j=1,2}\left|w_{+,j}^{\uparrow}\right|\sin\left(E_{j}t+\arg\left(w_{+,j}^{\uparrow}\right)\right),$ (11)
where we took into account the particle-hole symmetry relation $w_{-,j}^{\sigma}=(w_{+,j}^{\sigma})*$. It is important to emphasise that after calculating probability at a defect site, one finds that the two spin components exactly cancel oscillating parts of each other for a single value of $j$ due to the different phase of oscillations between $sin$ and $cos$ in equations (10),(11) above
$|\Psi_{\sigma,d}^{j}(t)|^{2}=2\left|w_{+,j}^{\uparrow}\right|^{2}$
×$\times\left(1+\text{sign}(\sigma)\cos\left(2E_{j}t+2\arg(w_{+,j}^{\uparrow})\right)\right),$ (12)
with $\text{sign}(\sigma)=\pm$. Thus, the contribution of each single pair of trapped states is constant and due to the initial condition of uniformly distributed state over entire system it could not describe the peak in probability.
Figure 4: The comparison of density evolution at searched defect node (blue solid line) and another distanced node without defect (gray dashed line) with time for a walker parameters ‘1’ and defect marked by $(5\pi/8,\pi/2)$. The red dash-dotted line corresponds to the evolution described by only two pairs of trapped states that maximize product of overlaps and given by Eq.(Constant search time algorithm via topological quantum walks). The evolution from trapped states approximates with very good precision the first few peaks in probability and the time of their occurence.
However, the crucial contribution comes from the superposition between two pairs of trapped states and is manifested in the last oscillating term in the total density evolution at the defect:
$\left|\Psi_{\uparrow,d}(t)\right|^{2}+\left|\Psi_{\downarrow,d}(t)\right|^{2}=4\left(\left|w_{+,1}^{\uparrow}\right|^{2}+\left|w_{+,2}^{\uparrow}\right|^{2}\right)+8$
×$\times\left|w_{+,1}^{\uparrow}\right|\left|w_{+,2}^{\uparrow}\right|\cos\left(\left(E_{1}-E_{2}\right)t+\arg w_{+,1}^{\uparrow}-\arg w_{+,2}^{\uparrow}\right).$ (13)
The period of this oscillation is different from the fast oscillations of each trapped states and defined by the energy difference between the positive energy levels of two pairs of trapped states $(E_{1}-E_{2})$. The plot of evolution of probability on defect in Fig.4 shows that such simple expression captures very well the structure of oscillations, the time of first peak appearance and the max probability on the first period on a defect. In Fig.4 the results are presented for a walker ‘1’, and the same is shown for a walker ‘2’ in Supplemental material. An another way to check the agreement of predicted period of oscillations with the actual oscillation of probability on a defect node we used numerical Fourier transform. The agreement with period defined by $(E_{1}-E_{2})$ has errors below $2\%$ on the timescale of evolution $T=5000$ for system with $\sqrt{N}=40$ [22]. Note that such definition of period through the difference of eigenstates of operator is different from the common one used in Grover’s search in QRW [23, 7]. Specifically, the period definition we use here is between states from different particle-hole symmetric pairs rather that between states from single pair.
The change of time asymptotic that is shown in Fig.2(b) happens when the system size becomes larger that diameters of both trapped state pairs. In Supplemental material [21] we show that the state from the pair with smaller overlap in Fig.2(c) has larger radius oflocalization ($95\%$ of the probability density). This pair of states determines the chance in the search time asymptotic as shown in Fig. 2b.
Finally, we note on stability of algorithm in comparison with Grover’s search. In Supplemental material [21] we present a comparison from Fig.2(d) for the disordered systems with disorder being implemented in rotation angles. In the large disorder cases walker ‘2’ shows similar or even better performance that Grover’s search implemented via QRW, and is also advantageous against classical search.
In conclusion, we demonstrated that a fast quantum search algorithm can be realized in topological split-step quantum random walks. We proposed a concrete implementation of this search consisting of the following steps, namely (i) Check the appearance of pair of trapped states near a defect and overlap product maximization criteria for a small system; (ii) apply the same encoding of searched site in the large system with the same walker parameters;(iii) initialize uniform state in the large system; (iv) the high density peak in time will reveal the searched node position [24]. Our algorithm admittedly relies on the proper tuning of walker and searched node parameters. At the same time this property also allows for a protection against disorder with $\Theta_{1}$ parameter values being not within efficient search interval. Interestingly, in the search algorithm we present here the ‘structure’ is induced in a quantum way (via dynamically appearing interference pattern due to overlap with trapped states) after fixed amount of time, i.e. it is not known a-priori where the pairs of trapped states are localized. In such formulation, our algorithm is suitable to make fast detectors of emergent defects in topological split-step quantum random walks. In addition, we point out that the dependence of algorithm of superposition of only four states makes it less sensitive to disorder rather that Grover’s algorithm, which performs optimally when a significant percentage of states contributed into constructive interference. Our search algorithm, as formulated in this work, can be immediately tested in photonic and synthetic lattice quantum walk experiments that realize topological QRW [25, 26, 27, 28].
Acknowledgements. We benefited from discussions with Anton Akhmerov, Yaroslav Herasymenko, Margarita Davydova and Carlo Beenakker. This project was supported by the Kavli Foundation. This publication is part of the project Engineered Topological Quantum Networks (with Project No. VI.Veni.212.278) of the research program NWO Talent Programme Veni Science domain 2021 which is financed by the Dutch Research Council (NWO). GJ acknowledges the research program “Materials for the Quantum Age” (QuMat) for financial support. This program (registration number 024.005.006) is part of the Gravitation program financed by the Dutch Ministry of Education, Culture and Science (OCW).
Supplemental material for “Constant search time algorithm via topological quantum walks”
by D. O. Oriekhov, Guliuxin Jin, and Eliska Greplova
Appendix S1 Properties of split-step quantum random walk
In this section we present a number of properties for split-step QRW that realizes topological phases in the case without defect nodes. Using Eq.(1) from the main text, and assuming the infinite system in both directions, we find the following equation for spectrum of translation-invariant unit cell:
$\lambda^{2}+2\lambda\cos\left(\Theta_{2}/2\right)\left(\sin^{2}\left(\Theta_{1}/2\right)\cos\left(2k_{y}\right)-\cos^{2}\left(\Theta_{1}/2\right)\cos\left(k_{x}\right)\right)+\lambda\sin\left(\Theta_{1}\right)\sin\left(\Theta_{2}/2\right)\left(\cos\left(k_{x}-2k_{y}\right)+1\right)+1=0.$ (S1)
The solution of this equation, $\lambda_{\pm}=e^{iE}$, gives a particle-hole symmetric spectrum of quasi-energies. The quasi-energy is given by
$E=\pm\arccos\frac{b}{\sqrt{2b^{2}-4}},$
$b=2\cos\left(\Theta_{2}/2\right)\left(\sin^{2}\left(\Theta_{1}/2\right)\cos\left(2k_{y}\right)-\cos^{2}\left(\Theta_{1}/2\right)\cos\left(k_{x}\right)\right)+\sin\left(\Theta_{1}\right)\sin\left(\Theta_{2}/2\right)\left(\cos\left(k_{x}-2k_{y}\right)+1\right)$ (S2)
The topological phases and corresponding Chern numbers were defined in Ref.[8]. The phase separation lines are characterized by the gap closing. For the technical purposed of identifying different types of search speed up discussed in the main text, we plot a map of gap value: the minimal gap value between two bands at $E=0$ or $E=\pm\pi$ (see Fig.S1(a) ) and the maximal gap value around the same energies. From such maps we find that the full gap closing happens only at lines $\Theta_{2}=2\pi n$ with $n$ being integer.
$(a)$Refer to caption $(b)$Refer to caption
Figure S1: The regions of one gap closing in panel (a) and both gaps closing in panel (b). Note the agreement between panel (a) and phase separation lines position.
The colormap of search dependence on defect parameters for the walker ‘1’ shown in Fig.2 of the main text is presented in Fig.S2. The dependence on $\Theta_{2}^{def}$ is absent, which motivates the studying of only different $\Theta_{1}^{def}$ values in the main text.
Figure S2: @D colormap of dependence on defect parameters for a walker ‘1’. The optimal search is performed on a line - which demonstrates that the dependence on $\Theta_{2}^{def}$ is absent.
Appendix S2 The comparison of product of overlaps with optimal search regions
In this section we analyze how precise the correspondence between maximization of $|<d|n><n|i>|^{2}$ and $(\sum_{n=max_{1},max_{2}}|<d|n><n|i>|)^{2}$ is compared to the regions of optimal search in parameter space for a fixed defect parameters. For this purpose we compare the max search probability curves plotted along two lines depicted as dashed cyan in the first panel of Fig.S3. One of these lines is a phase separation line and second one is parallel to phase separation and contains a walker `$1^{\prime}$ discussed in the main text. This second line crosses different topological phases. In the second and third panels of Fig.S3 we plot the comparison of max probability found on time interval $T=1000$ for the system size with $\sqrt{N}=40$ with values of different overlaps. The second and third panels in Fig.S3 clearly demonstrate that the combination of maximum and second maximum products of overlaps - the contribution of two pairs of trapped states - describes the probability peak most precisely. This leads to a conclusion that efficient search is performed by two pairs of trapped states in this system in all cases studied.
Figure S3: Comparison of maximal product of overlaps described in the main text found over all eigenstates of evolution operator $\bar{U}$ for a two lines in parameter space plotted as dashed cyan lines in the first panel. The second panel corresponds to a line going through walker `$1^{\prime}$ and the third panel corresponds to a phase separation line.
Next, we analyze the structure of trapped states and their position in spectrum for the example walker parameters labeled as `$2^{\prime}$. The analysis is similar to what is done in the main text and is presented in three panels of Fig.S4: comparison of search probability with overlap criteria in the first panel, the set of overlaps for each eigenstate of $\bar{U}$ in second panel and the comparison of probability evolution at the defect with the one defined by the two pairs of the trapped states in the third panel. However, in this case we find that the two pairs of trapped states that mainly contribute to the efficient search are placed near zero energy. Also there is no gap in spectrum around this energy, but the in-gap trapped states placed around $E=\pm\pi$ do not contribute into the constructive interference relevant for successful search.
Figure S4: (a) The comparison of search probability with overlap criteria for walker ‘2’. (b) The identification of states that mainly contribute into constructive superposition on a searched node for walker ‘2’ parameters. (c) The comparison of density evolution at searched defect node (blue solid line) with the superposition (red dashed line) of trapped states identified from panel (b).
Appendix S3 Structure of trapped states in coordinate space
In this section we analyze how extended are the trapped states that mainly contribute to the search algorithm. From the search time dependence on system size shown in Fig.2(b) in the main text, we conclude that the more extended trapped state has radius $R\approx 50$ for walker ‘1’ and $R\approx 100$ for walker ‘2’ and defect parameter $\Theta_{1}^{def}=5\pi/8.$. The radius of trapped state was extimated as twice smaller that the system size where transition in time asymptotic happens. To further analyze the trapped state radius dependence on system size, we plot it in Fig.S5. The set of tested system sizes covers the region where the trapped states are having overlap with itself due to periodic boundary condition. To reduce the numerical precision issue related to identifying density of quickly decaying state far from its center, we plot a dependence of radius which covers $80\%$ and $95\%$ of density. Two plots show results for a state taken from each pair of trapped states discussed in the main text in Fig.3(b). The states from second pair have larger radius and it stronger depends on system size. Thus, the change in search time asymptotic is mostly defined by this second pair of trapped states, that has second maximal overlap value.
Figure S5: Radius of trapped state as a function of system size. Left panel corresponds to a trapped state with largest overlap $|\langle d|n\rangle\langle n|i\rangle$ in Fig.3(b) of the main text, and right panel corresponds to the state with second largest overlap.
Appendix S4 Stability of search algorithm with disorder
In this section we make a stability comparison of the proposed algorithm of search based on trapped states with Grover’s algorithm in split-step quantum random walk described in Ref.[7]. The size of the systems is taken to be the same - $512\times 512$ nodes, well above the saturation time threshold for both example walker parameters described in the main text. Since the Grover’s algorithm relies on the coherent superposition of most states present in the system, one could expect that disorder break such constructive superposition faster that for a few trapped states. To check that, we implement the angle disorder of spin rotation matrices on each node. For equation (2) in the main text such disorder is introduced as
$R(\Theta)\to R(\Theta+\delta\Theta_{xy}),\quad\delta\Theta\in[-\delta\Theta_{max},\delta\Theta_{max}],$ (S3)
where the distribution of $\delta\Theta_{xy}$ is uniform in the interval $[-\delta\Theta_{max},\delta\Theta_{max}]$. Similarly, for the split-step QRW analyzed in Ref.[7] and used here as Grover’s algorithm benchmark, the disorder is implemented as
$C_{x}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&i\\ i&1\end{pmatrix}\to C_{x}(\delta\Theta)=\frac{1}{\sqrt{2}}\begin{pmatrix}1&i\exp(i\delta\Theta)\\ i\exp(-i\delta\Theta)&1\end{pmatrix},$ (S4)
$C_{y}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-i\\ -i&1\end{pmatrix}\to C_{y}(\delta\Theta)=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-i\exp(i\delta\Theta)\\ -i\exp(-i\delta\Theta)&1\end{pmatrix},$ (S5)
The results of simulation for equal intervals of disorder angle are presented in Fig.S6.
The role of disorder in reducing efficiency of search algorithms in visible in all plots. However, for one of the two walkers, the search probability is always staying higher that the classical probability on the first period of oscillation. Notably, the probability of Grover’s search drops dramatically. In the presented example of disorder configuration and type, the trapped state are demonstrating relatively stable behavior as a basis of search algorithm, even outperforming Grover’s example for large disorder. Still, we point out that the more systematic study should be made to define the best performance conditions of different suboptimal quantum algorithms and compared with Grover’s one in noisy quantum setups. Also one should note that these algorithms still outperform classical search, which behaves as $P_{classical}(t)=t/N$ with $t$ being number of check ups of individual nodes.
Figure S6: The search probability calculated for 2 walker parameters analyzed in the main text (see Fig.2) and compared with Grover’s algorithm in split-step QRW from Ref.[7] and with classical search with linearly increasing probability. The system size is $512\times 512$. Disorder strength varies from small in panel (a) $\delta\Theta_{max}=0.05$, to (b) $\delta\Theta_{max}=0.15$, (c) $\delta\Theta_{max}=0.35$, and large in panel (d) $\delta\Theta_{max}=0.5$. The walker $2$ shows best performance at large disorder out of all compared search procedures.

### v3

@@
-Constant search time algorithm via topological quantum walks
+Dynamical localization in 2D topological quantum random walks
 Abstract
-It is well-known that quantum algorithms such as Grover’s can provide a quadradic speed-up for unstructured search problems.
-By adding topological structure to a search problem, we show that it is possible to achieve a constant search-time quantum algorithm with a constant improvement of the search probability over classical search.
-Specifically, we study the spatial search algorithm implemented by a two-dimensional split-step quantum random walks that realize topologically nontrivial phases and show the asymptotic search behavior is constant with growing system size.
-Using analytical and numerical calculations, we determine the efficient search regions in the parameter space of the quantum walker.
-These regions correspond to pairs of trapped states formed near a lattice defect.
-By studying the spectral properties of the discrete time-evolution-operators, we show that these trapped states have large overlap with the initial state.
-This correspondence, which is analogous to localization by constructive interference of bound states, makes it possible to reach the best possible search-time asymptotic and produce a disorder-protected fast search in quantum random walks.
-Quantum search algorithm, Grover’s search [1], is known as an algorithm that demonstrates quantum advantage over its classical analogue.
-Grover’s algorithm performs a search of marked element in an unstructured database with a given probability with asymptotic time being a square root of database element count.
-However, its implementation depends on an oracle operator - a way of distinguishing marking element that applies to a whole quantum state.
-The quantum random walks are known as an example of natural implementation of such an oracle operator [2, 3, 4, 5, 6].
-The discrete-time quantum random walks are defined as a set of shift and spin flip operators applied at the whole system at each time step.
-The successful search event is defined as a time of maximum localization of quantum walker on a defect when starting from a uniform distribution over all system.
-Although, different kinds of quantum walks demonstrate a Grover-type convergence of a spatial search in two and higher dimensions, the dependence on spin flip operator and position of defect site relative to remaining system is crucial and in many cases results in a decreased efficiency of search algorithm [5, 4, 7].
-The latter fact also serves as a motivation for the search for sub-optimal yet more robust to disorder and element positioning algorithms, that would provide quantum advantage.
-In the present paper we study the two-dimensional split-step quantum random walks that are analogues of Chern topological insulators [8] in crystalline solids.
-The similar split-step QRW in 1D were studied in Refs.[9, 10] and a number of our conclusions made below might work also in those cases.
-By analyzing the parameter spaces of quantum walker and defect, we find that there exist sets of parameters that demonstrate efficient search of a given defect (see a search probability map in Fig.1).
-Below we show that these regions in parameter space can be categorized into two distinct classes based on their position inside the topological phase or on the phase separation lines.
-Notably, the search time for the parameters inside both classes of regions demonstrates saturation to a constant search time with growing system size.
-The saturation of search time happens at different system sizes, while before that the square root time growth behavior takes place.
-Figure 1: The success probability of search algorithm as a function of QRW parameters $\Theta_{1,2}$ (see Eq.(2)) for a defect with $\Theta_{1}^{def}=5\pi/8$, $\Theta_{2}^{def}=\pi/2$.
-High peaks are of order of magnitude larger than average localization probability and correspond to parameters for which quantum walker performs search efficiently.
+We study the dynamical localization of discrete time evolution of topological split-step quantum random walk (QRW) on a single-site defect starting from a uniform distribution.
+Using analytical and numerical calculations, we determine the high localization probability regions in the parameter space of the quantum walker.
+These regions contain two or more pairs of trapped states, forming near a lattice defect.
+By investigating the spectral properties of the discrete-time evolution operators, we show that these trapped states have large overlap with the initial uniformly distributed state, thus offering a simple interpretation of the localization effect.
+As this localization scheme could be interpreted as a variation of spatial quantum search algorithm, we compare the localization probability and time with other types of two-dimensional quantum walks that do not have topological phases and realize localization time scaling similar to Grover’s algorithm.
+Finally we show that mechanism of localization we identified is robust against the influence of disorder.
+I Introduction
+The quantum random walks were introduced in Ref.
+[1] as a natural quantum version of classical random walks.
+Due to interference effects and the state superposition, the average travel path of walkers is much longer than their classical counterpart.
+The clear structure of possible physical systems that could realize the quantum walks [1] attracted much attention.
+In particular, it was shown that quantum walks are a possible way to implement Grover’s algorithm [2] in the case of absence of high-quality multi-qubit quantum computer [3, 4, 5, 6, 7, 8].
+It was also shown that the so-called Grover’s oracle operator has a natural realization in quantum walks and is related to dynamical localization phenomena [3, 4, 5, 6, 7, 8].
+In particular, the discrete-time quantum random walks are defined as a set of shift and coin-flip operators applied to the whole system at each time step [5, 9, 1, 3].
+The successful search event is defined as a time of maximum localization of quantum walker on a defect when starting from a uniform distribution over the whole system.
+Going forward, we will refer to the the time of appearance of the first maximum in localization probability as search time.
+Figure 1: The localization probability of topological quantum walk as a function of spin rotation parameters $\Theta_{1,2}$ (see Eq.
+(2)) for a defect with $\Theta_{1}^{def}=5\pi/8$, $\Theta_{2}^{def}=\pi/2$.
+High peaks are of two orders of magnitude larger than average localization probability and correspond to parameters for which quantum walker performs search efficiently.
 Inset: geometry of split-step QRW on a triangular lattice with three shift vectors ${\boldsymbol{v}}_{i}$ solid ones correspond to shift of spin up components, dashed $-{\boldsymbol{v}}_{i}$ for spin down.
-The constant search time asymptotic was previously indentified only for electric Dirac QRW with non-local marking of defect in the form of Coulomb potential [11, 12].
-The non-locality creates an important difference from our algorithm.
-Namely, here the constant search time is achieved with local marking, thus making it directly implementable in experimental memory setup.
-Below we analyze in detail the time evolution of states for a range of quantum walker parameters and study the overlap of initial state with the eigenstates of a unitary operator corresponding to a single time step of QRW.
-We demonstrate that the search speedup with constant time asymptotic is related to the appearance of trapped states of a specific structure near the defect: the two pairs of trapped states should simultaneously have a large matrix element of coordinate operator of the defect and the large overlap with initial uniformly distributed state.
-In the mathematical sense the trapped states are defined as a states that have compact support and zero density elsewhere.
-Their analogue in quantum mechanics are the strongly localized bound states on impurity in a system.
+In the most common mechanism of localization, at least several eigenstates of QRW evolution operator create a constructive superposition on a defect at a certain time [6, 5].
+Little is known about the precise values of optimal parameters of defect that maximize localization relative to the surrounding system.
+In addition, it was shown that certain positioning of defects could significantly decrease localization probability [10].
+Here, we study the two-dimensional split-step quantum random walks analogous to Chern type topological insulators in crystalline solids [11].
+For such a quantum walk, it could be expected that topology creates a new mechanism of localization, similar to in-gap states in solids.
+Following this motivation we explore how topology affect the localization mechanism.
+The topological split-step QRWs in 1D space were studied in Refs.‘[12, 9].
+Here, we analyze the parameter spaces of the quantum walker and the defect.
+We find that there exist sets of parameters that deterministically demonstrate a high localization probability at a given defect (see a localization probability map in Fig.
+1).
+In Sec.
+II we show that these regions in parameter space can be categorized into two distinct classes based on their position inside the topological phase or on the phase separation lines.
+Notably, we find the search time for the parameters inside both classes of regions demonstrates the saturation to a constant with growing system size.
+The saturation of search time happens at different system sizes, while before this saturation the square root time growth behavior takes place.
+The localization after constant time, regardless of system size, was previously identified only for electric Dirac QRW with non-local marking of defect in the form of Coulomb potential [13, 14].
+The non-locality of defects creates an important difference from our scenario: our protocol does not require a large number of nodes to form a defect structure.
+In the algorithmic terminology, in our work the constant search time is achieved with local marking comparing to non-local extended marking of Coulomb potential.
+In Sec.
+III we analyze the time evolution of QRW states for a range of parameters in detail, and study the overlap of initial state with the eigenstates of the unitary operator corresponding to a single time step of QRW.
+Particularly, we demonstrate that the search speedup with constant time asymptotic is related to the appearance of trapped states of a specific structure near the defect: the two pairs of trapped states should simultaneously have a large matrix element with coordinate operator of the defect, and have large overlap with uniformly distributed initial state.
+In mathematical terms the trapped states are defined as a states that have compact support and zero density elsewhere.
+Their analogue in condensed matter are the strongly localized bound states on impurity in a system.
 We note that a number of mathematically rigorous studies have been made for spectral properties of quantum random walks and trapped states [6, 15, 12, 16, 17, 18] together with their dynamics [19, 20].
 However, the possible efficiency of search algorithm based on exploiting the particular structure of QRW was not previously addressed.
-The strong dependence of the trapped states structure on defect and walker parameters also suggests a scenario of protecting quantum search from a random disorder by an appropriate fine-tuning of marking parameters for the searched element.
-The quantum random walk deployed for algorithmic speed-up in this work was firstly introduced in Ref.[8] as a way to simulate topological insulators. Its single time step
+The strong dependence of the trapped states structure on defect and walker parameters also suggests a scenario of protecting quantum search from a random disorder by an appropriate fine-tuning of marking parameters for the searched defect element.
+The corresponding study is presented in Sec.
+IV and the concluding remarks are given in Sec.
+V.
+II The QRW model and localization phenomenon
+The quantum random walk analyzed in this work was first introduced in Ref.[11] as a way to simulate topological insulators. Its single time step
 $\Psi(t+1)=U(\Theta_{1},\Theta_{2})\Psi(t)$ (1)
-is defined via three rotations of spin and subsequent translations:
-$U(\Theta_{1},\Theta_{2})=T_{{\boldsymbol{v}}_{3}}R\left(\theta_{1}\right)T_{{\boldsymbol{v}}_{2}}R\left(\theta_{2}\right)T_{{\boldsymbol{v}}_{3}}R\left(\theta_{1}\right),$ (2)
+is defined via three rotations of the spin and subsequent translations:
+$U(\Theta_{1},\Theta_{2})=T_{{\boldsymbol{v}}_{3}}R\left(\theta_{1}\right)T_{{\boldsymbol{v}}_{2}}R\left(\theta_{2}\right)T_{{\boldsymbol{v}}_{1}}R\left(\theta_{1}\right),$ (2)
 $R(\theta)=\left[\begin{array}[]{cc}\cos(\theta/2)&-\sin(\theta/2)\\ \sin(\theta/2)&\cos(\theta/2)\end{array}\right].$ (5)
@@
 (b) Search time measured as a position of first probability peak as a function of system size for walker parameters indicated by crosses ‘1’ and ‘2’ in panel (a).
-(c) Search probability dependence on system size.
-The two regions of different dependence are present where probability follows Grover-type scaling $P\sim 1/\log(N)$ until the time asymptotic changes to constant in (b) and then switches to system size scaling $P\sim 1/N$ but having few orders of magnitude larger value than average density in the system.
-(d) Comparison of performance with Grover’s algorithm and repeated classical search for system size $\sqrt{N}=512$.
+(c) Localization probability dependence on system size.
+The two regions of different dependence are present where probability follows ‘Grover-type’ scaling $P\sim 1/\log(N)$ as non-topological QRW from Ref.[10] until the time asymptotic changes to constant in (b) and then switches to system size scaling $P\sim 1/N$ but having few orders of magnitude larger value than average density in the system.
+(d) Comparison of time-dependent localization probability with non-topological QRW [10] and repeated classical search for system size $\sqrt{N}=512$.
 Walker $2$ shows advantage over classical.
-In our setup the search algorithm goal is defined as a problem of finding a single node on a finite 2D lattice of size $\sqrt{N}\times\sqrt{N}$ unit cells starting from initial state $|i\rangle$.
+Note the scaling factor of $31e-$ 3 in the panel (d) for localization probabilities.
+In our setup the localization phenomena is expected to occur after certain amount of time steps at a single defect node on a finite 2D lattice of size $\sqrt{N}\times\sqrt{N}$ unit cells starting from initial state $|i\rangle$.
 We apply the periodic boundary condition in order to avoid the appearance of edge states.
-The searched node is marked by parameters $\Theta_{1,2}^{def}$ that are different from surrounding $\Theta_{1,2}$.
-To estimate the possibility of given quantum walker to perform search efficiently, we firstly scan the entire parameter range $\Theta_{1,2}$ for a given defect values $\Theta_{1,2}^{def}$.
-The performance of search algorithm is measured by a search probability.
-It is defined as a peak $\max_{t\in[0,T]}P_{def}(t)$ in time-dependent density of the system state at the defect node,
+The defect node is marked by parameters $\Theta_{1,2}^{def}$ that are different from surrounding $\Theta_{1,2}$.
+To estimate the possibility of given quantum walker to localize efficiently with high probability, we firstly scan the entire parameter range $\Theta_{1,2}$ for a given defect values $\Theta_{1,2}^{def}$.
+The measure of localization probability that is relevant for quantum search applications is defined as a peak $\max_{t\in[0,T]}P_{def}(t)$ in time-dependent density of the system state at the defect node,
 $P_{def}(t)=|\langle d|\Psi(t)\rangle|^{2},$ (6)
 where $|d\rangle=\sum_{\sigma=\uparrow,\downarrow}|x=def,\sigma\rangle$ is the defect coordinate operator with both components of spin $\sigma$, and $|i\rangle$ is the initial state of the system defined as $|i\rangle=\frac{1}{\sqrt{2N}}\sum_{\sigma,x}|x,\sigma\rangle$ with summation over entire system.
-The plot of the maximum value of density $P_{def}(t)$ on defect over the interval of time $T=1000$ for test system size $N=40\times 40$ sites is shown in Fig.1 for $\Theta_{1,2}^{def}=(5\pi/8,\pi/2)$.
-In Fig.2(a) we show the colormap representation of the same search probability distribution on the top of phase diagram in parameter space (see also Supplement [21] for other defect parameters).
-The islands of efficient search in parameter space fill a small areas placed in two distinct types of regions: inside topological phases and on phase separation lines.
-We compare the search time and probability behavior with growing system size for the walker parameters marked by ‘1’ and ‘2’ in the Fig.2(a), that realize best search probability within its regions.
-Such results are shown in Fig.2(b).
-With the growing system size the behavior of search time defined via first probability peak demonstrates qualitative change - from $T\sim\sqrt{N}$ dependence to $T\sim const$ in both curves.
-As we noted previously, such speed-up was observed only for the Dirac QRW with particular type of Coulomb-like disorder potential [11, 12].
-In Fig.2(c) the search probability dependence of walkers ‘1’ and ‘2’ on system size is compared with average density at non-defect nodes and Grover’s algorithm search probability.
+The plot of the maximum value of density $P_{def}(t)$ on defect over the interval of time $T=1000$ for test system size $N=40\times 40$ sites is shown in Fig.
+1 for $\Theta_{1,2}^{def}=(5\pi/8,\pi/2)$.
+In Fig.
+2(a) we show the color map representation of the same localization probability distribution on the top of phase diagram in parameter space.
+For other positions of defect $\Theta_{1,2}^{def}$ in parameter space the colormaps of localization probability are shown in Supplemental material [23].
+The islands of efficient localization in parameter space fill a small areas placed in two distinct types of regions: inside topological phases and on phase separation lines.
+We compare the search time and localization probability behavior with growing system size for the walker parameters marked by ‘1’ and ‘2’ in the Fig.2(a), that realize maximal values of probability within its regions.
+In Fig.
+2(b), the search time is plotted against the square root of the system for the walkers ’1’ and ’2’.
+With the growing system size the behavior of search time demonstrates qualitative change - from $T\sim\sqrt{N}$ dependence to $T\sim\textit{const.}$.
+for both walkers.
+As we noted previously, such such a dependence of localization maximum on time was observed only for the Dirac QRW with particular type of Coulomb-like disorder potential [13, 14].
+In Fig.
+2(c) the localization probability of walkers ‘1’ and ‘2’ is compared to average density at non-defect nodes.
 We find that the constant search time regime corresponds to the $P_{def}(t)=O(1/N)$ with a large numerical prefactor.
-This decrease of search probability with system size $N$ is faster than Grover’s one, $P_{def}^{Grover}=O(1/log\,N)$.
-The search probability dependence in the form $O(1/N)$ arises due to the trapped states having finite support.
-The system size grows beyond the trapped state support size corresponds to saturation of search time asymptotic to constant and clear change in probability dependence curve in Fig.2(c).
-To check the dependence of search probability on defect parameters for given walker we performed the same 2D parameter space calculation taking walker "$1"$ and checking $\Theta_{1,2}^{def}\in[-2\pi,2\pi]$.
-It is found that the search probability does not depend on $\Theta_{2}^{def}$ [21], thus the more informative is to present a dependence on $\Theta_{1}^{def}$ as shown in Fig.3(a).
-The asymptotic scaling of $O(1/N)$ in search probability does not reach the optimal Grover scaling, but still provides constant improvement over the classical search algorithm.
-The probability of classical algorithm grows as $t/N$ with multiple trials, $t$ (one per time step).
-For this purpose, in Fig.2(d) we present a comparison of probability on searched node as a function of time for two walkers marked in Fig.2(a), Grover’s search for split-step QRW from Ref.[7] and a classical search.
-The walker ‘2’ demonstrates twice as large probability as classical search, while the walker ‘1’ follows the classical search probability for approximately 100 steps.
-In terms of detection via measurements, higher probability on a node means that the detector resolution used to identify position could be lower - in particular, for walker ‘1‘ the detection resolution could be $100$ times smaller that for classical probes of individual nodes.
-The Grover’s algorithm performance demonstrates even higher probability at the same time scale, which makes it advantageous in the non-disordered system.
-As we point out below, that advantage of Grover’s search might be lost in a disordered systems.
-To describe the origin of different speed-ups in this system and why the regions of efficient search represent a small fraction of parameter space, we analyze the spectrum and localization properties of eigenvectors of a unitary operator of system with one defect.
+For the algorithmic comparison it is important to note that this decrease of localization probability with system size $N$ is faster than for non-topological QRW discussed in Ref.
+[10] $P_{def}^{Grover}=O(1/log\,N)$.
+The localization probability dependence in the form $O(1/N)$ arises due to the finite support of the trapped states.
+When the system size grows beyond the support of the trapped state, the search time asymptotic get saturated and we observe a clear dependence in the localization probability on the system size, see Fig.
+2(c).
+For the fixed parameters of quantum walker ‘1’ we visualized the distribution of localization probability in the space of defect parameters $\Theta_{1,2}^{def}\in[-2\pi,2\pi]$.
+The corresponding discussion in Appendix A shows that the localization probability depends only on $\Theta_{1}^{def}$ parameter, and is constant along the $\Theta_{2}^{def}$.
+Thus, it is more informative to present a dependence on $\Theta_{1}^{def}$ as shown in Fig.3(a) to compare with overlap criteria introduced below.
+[…]
+III Contribution of trapped states to dynamical localization
+To describe the origin of dramatically different localization behaviors for different parameters of the topological QRW and to explain why the regions of efficient localization represent only a small fraction of the parameter space, we analyze the spectrum and localization properties of eigenvectors of a unitary operator of the system with one defect.
 The probability evolution at a defect site at discrete times $t$ can be rewritten through the unitary operator of QRW,
@@
 Substituting the eigenbasis decomposition of the $\bar{U}$ operator, $\bar{U}|n,\pm\rangle=e^{i\pm E_{n}}|n,\pm\rangle$, we have
-$P_{def}(t)=\bigg{|}\langle d|\sum_{n,\lambda=\pm}e^{i\lambda E_{n}t}|n,\lambda\rangle\langle n,\lambda|i\rangle\bigg{|}^{2}$ (8)
+$P_{def}(t)=\bigg{|}\langle d|\sum_{n,\lambda=\pm}e^{i\lambda E_{n}t}|n,\lambda\rangle\langle n,\lambda|i\rangle\bigg{|}^{2}.$ (8)
 The saturation of search time to constant scaling after system size becomes large enough suggests that the properties of states that mainly contribute to the Eq.(8) do not change with system size.
@@
 [6, 15] in the context of QRWs staying localized around its initial starting node over infinite evolution time.
-We investigate the contribution of different eigenstates of $\bar{U}$ to the evolution of probability, $P(t)$, in Fig.3(b) by comparing the product of overlaps $|\langle d|n,\lambda\rangle\langle n,\lambda|i\rangle|^{2}$.
-The panels in Fig.3(b) correspond to different parameters of defect denoted by vertical dashed lines for walker `$1^{\prime}$ in Fig.3(a).
-Figure 3: (a) Comparison of maximum value of probability on defect for a walker ‘1’ from Fig.2 as a function of defect parameter $\Theta_{1}^{def}$ with the maximal value of squared of overlaps defined in Eq.(9) with $\mathcal{M}$ containing one and two states respectively.
+We investigate the contribution of different eigenstates of $\bar{U}$ to the evolution of probability, $P(t)$, in Fig.
+3(b) by comparing the product of overlaps $|\langle d|n,\lambda\rangle\langle n,\lambda|i\rangle|^{2}$.
+The panels in Fig.
+3(b) correspond to different parameters of defect (listed above each panel) denoted by vertical dashed lines for walker `$1^{\prime}$ in Fig.
+3(a).
+Figure 3: (a) Comparison of maximum value of probability on defect for a walker ‘1’ normalized by numerically achieved maximum on the interval from Fig.2 as a function of defect parameter $\Theta_{1}^{def}$ with the maximal value of squared sum of overlaps defined in Eq.(9) with $\mathcal{M}$ containing one and two states respectively.
 All quantities are normalized to scale (0,1) for the purpose of comparison.
@@
 Trapped states differ from other states by significantly higher value of overlap.
-The scale of $x$-axis in each panel of Fig.3(b) shows that defect parameters with higher search probability also correspond to higher overlap product for a few states - in this case two pairs of particle-hole symmetric states.
+The scale of $x$-axis in each panel of Fig.
+3(b) shows that defect parameters with higher localization probability also correspond to higher overlap product for several states - in this case two pairs of particle-hole symmetric states.
 To further verify this observation, we compare the maximal value of overlap product
-$W_{\mathcal{M}}=|\sum_{n\in\mathcal{M},\lambda=\partial}\langle d|n,\lambda\rangle\langle n,\lambda|i\rangle|^{2}$ (9)
+$W_{\mathcal{M}}=\left|\sum_{n\in\mathcal{M},\lambda=\partial}\langle d|n,\lambda\rangle\langle n,\lambda|i\rangle\right|^{2}$ (9)
 of a single eigenstate of $\bar{U}$ with the normalized probability dependence $max[P_{def}(t)]$ as function of $\Theta_{1}^{def}$.
 The comparison is made by taking a single pair of states that maximize overlap with set of indices =$\mathcal{M}=\{n_{max}\}$, as well as two pairs of states =$\mathcal{M}=\{n_{max},n_{second\,max}\}$.
-The results presented in Fig.3(a) indicate that while =$\mathcal{M}=\{n_{max}\}$ selection overestimates the region of good search, the =$\mathcal{M}=\{n_{max},n_{second\,max}\}$ very precisely describes the high search probability peak.
-To further check this correspondence between the evolution of described by two pairs of states with maximal overlap and the max probability distribution, we plotted the same comparison for a defect $\Theta_{1,2}^{def}=(5\pi/8,\pi/2)$ along phase separation line and the line with walker `$1^{\prime}$ in Supplemental material.
-The correspondence for a test system size with $\sqrt{N}=40$ shows that efficient search for such split-step QRW is mostly determined by the appearance of two pairs of trapped states having large overlap with both defect and initial state.
+The results presented in Fig.
+3(a) indicate that while =$\mathcal{M}=\{n_{max}\}$ selection overestimates the region of good search, the =$\mathcal{M}=\{n_{max},n_{second\,max}\}$ very precisely describes the high search probability peak.
+To further check this correspondence between the evolution of described by two pairs of states with maximal overlap and the maximum probability distribution, we show the same comparison for a defect $\Theta_{1,2}^{def}=(5\pi/8,\pi/2)$ along phase separation line and the line with walker `$1^{\prime}$ in Appendix B.
+The correspondence for a test system size with $\sqrt{N}=40$ shows that efficient search for such split-step QRW is mostly determined by the appearance of two pairs of trapped states that have large overlap with both defect and initial state.
 Finally, we perform a comparison of time evolution defined by two pairs of trapped states that maximize overlap criteria with the evolution of probability on the defect.
@@
 where we took into account the particle-hole symmetry relation $w_{-,j}^{\sigma}=(w_{+,j}^{\sigma})*$.
-It is important to emphasise that after calculating probability at a defect site, one finds that the two spin components exactly cancel oscillating parts of each other for a single value of $j$ due to the different phase of oscillations between $sin$ and $cos$ in equations (10),(11) above
+It is important to emphasize that after calculating probability at a defect site, one finds that the two spin components exactly cancel oscillating parts of each other for a single value of $j$.
+That happens due to the different phases of oscillations between $sin$ and $cos$ in Eqs.(10) and (11) above, and results in:
 $|\Psi_{\sigma,d}^{j}(t)|^{2}=2\left|w_{+,j}^{\uparrow}\right|^{2}$
@@
 with $\text{sign}(\sigma)=\pm$.
-Thus, the contribution of each single pair of trapped states is constant and due to the initial condition of uniformly distributed state over entire system it could not describe the peak in probability.
+Thus, the contribution of each single pair of trapped states is constant and due to the initial condition of uniformly distributed state over entire system it could not describe the high peak in localization probability.
 Figure 4: The comparison of density evolution at searched defect node (blue solid line) and another distanced node without defect (gray dashed line) with time for a walker parameters ‘1’ and defect marked by $(5\pi/8,\pi/2)$.
-The red dash-dotted line corresponds to the evolution described by only two pairs of trapped states that maximize product of overlaps and given by Eq.(Constant search time algorithm via topological quantum walks).
-The evolution from trapped states approximates with very good precision the first few peaks in probability and the time of their occurence.
+The red dash-dotted line corresponds to the evolution described by only two pairs of trapped states that maximize product of overlaps and given by Eq.
+(III).
+The evolution from trapped states approximates with very good precision the first few peaks in probability and the time of their occurrence.
 However, the crucial contribution comes from the superposition between two pairs of trapped states and is manifested in the last oscillating term in the total density evolution at the defect:
@@
 ×$\times\left|w_{+,1}^{\uparrow}\right|\left|w_{+,2}^{\uparrow}\right|\cos\left(\left(E_{1}-E_{2}\right)t+\arg w_{+,1}^{\uparrow}-\arg w_{+,2}^{\uparrow}\right).$ (13)
-The period of this oscillation is different from the fast oscillations of each trapped states and defined by the energy difference between the positive energy levels of two pairs of trapped states $(E_{1}-E_{2})$.
-The plot of evolution of probability on defect in Fig.4 shows that such simple expression captures very well the structure of oscillations, the time of first peak appearance and the max probability on the first period on a defect.
-In Fig.4 the results are presented for a walker ‘1’, and the same is shown for a walker ‘2’ in Supplemental material.
-An another way to check the agreement of predicted period of oscillations with the actual oscillation of probability on a defect node we used numerical Fourier transform.
-The agreement with period defined by $(E_{1}-E_{2})$ has errors below $2\%$ on the timescale of evolution $T=5000$ for system with $\sqrt{N}=40$ [22].
-Note that such definition of period through the difference of eigenstates of operator is different from the common one used in Grover’s search in QRW [23, 7].
-Specifically, the period definition we use here is between states from different particle-hole symmetric pairs rather that between states from single pair.
-The change of time asymptotic that is shown in Fig.2(b) happens when the system size becomes larger that diameters of both trapped state pairs.
-In Supplemental material [21] we show that the state from the pair with smaller overlap in Fig.2(c) has larger radius oflocalization ($95\%$ of the probability density).
+The period of this oscillation is different from the fast oscillations of each pair of trapped states and is defined by the energy difference between the positive energy levels of two pairs of trapped states $(E_{1}-E_{2})$.
+Fig.
+4 shows that expression in Eq.
+(III) captures very well the structure of oscillations of localization probability despite its simplicity.
+In Fig.
+4 the results are presented for a walker ‘1’.
+We show the equivalent result walker ‘2’ in Appendix B.
+To further verify the agreement of predicted period of oscillations with the actual oscillation of probability on a defect node we used numerical Fourier transform.
+The agreement with period defined by $(E_{1}-E_{2})$ has errors below $2\%$ on the timescale of evolution $T=5000$ for system with $\sqrt{N}=40$.
+It was observed [24, 10] that in non-topological QRWs, that realize localization mechanism similar to Grover’s algorithm, and, the period of oscillation is defined by an energy difference within the same particle-hole symmetric pair of states at lowest energies.
+The difference between two definitions of the period shows that the mechanism of localization in topological QRWs is different comparing to non-topological QRWs studied in literature.
+The change of time asymptotic that is shown in Fig.
+2(b) happens when the system size becomes larger that diameters of both trapped state pairs.
+In Appendix C we show that the state from the pair with smaller overlap in Fig.2(c) has larger radius of localization ($95\%$ of the probability density).
 This pair of states determines the chance in the search time asymptotic as shown in Fig.
 2b.
-Finally, we note on stability of algorithm in comparison with Grover’s search.
-In Supplemental material [21] we present a comparison from Fig.2(d) for the disordered systems with disorder being implemented in rotation angles.
-In the large disorder cases walker ‘2’ shows similar or even better performance that Grover’s search implemented via QRW, and is also advantageous against classical search.
-In conclusion, we demonstrated that a fast quantum search algorithm can be realized in topological split-step quantum random walks.
-We proposed a concrete implementation of this search consisting of the following steps, namely (i) Check the appearance of pair of trapped states near a defect and overlap product maximization criteria for a small system; (ii) apply the same encoding of searched site in the large system with the same walker parameters;(iii) initialize uniform state in the large system; (iv) the high density peak in time will reveal the searched node position [24].
-Our algorithm admittedly relies on the proper tuning of walker and searched node parameters.
-At the same time this property also allows for a protection against disorder with $\Theta_{1}$ parameter values being not within efficient search interval.
-Interestingly, in the search algorithm we present here the ‘structure’ is induced in a quantum way (via dynamically appearing interference pattern due to overlap with trapped states) after fixed amount of time, i.e.
-it is not known a-priori where the pairs of trapped states are localized.
-In such formulation, our algorithm is suitable to make fast detectors of emergent defects in topological split-step quantum random walks.
-In addition, we point out that the dependence of algorithm of superposition of only four states makes it less sensitive to disorder rather that Grover’s algorithm, which performs optimally when a significant percentage of states contributed into constructive interference.
-Our search algorithm, as formulated in this work, can be immediately tested in photonic and synthetic lattice quantum walk experiments that realize topological QRW [25, 26, 27, 28].
+Figure 5: The localization probability calculated for 2 walker parameters analyzed in the main text (see Fig.2) and compared with the localization probability for QRW from Ref.[10] and with classical search with linearly increasing probability.
+The system size is $512\times 512$.
+Disorder strength varies from small in panel (a) $\Theta_{dis}=0.05$, to (b) $\Theta_{dis}=0.15$, (c) $\Theta_{dis}=0.35$, and large in panel (d) $\Theta_{dis}=0.5$.
+The walker $2$ shows best performance at large disorder out of all compared search procedures.
+IV Dynamical localization on defect in the presence of disorder
+Figure 6: The dynamical localization phenomena on a single-node defect averaged over $20$ configurations of disorder.
+The averaged localization in non-topological split-step QRW from Ref.[10] has a higher density of defect node rather than of a topological QRW due to a larger amount of states contributing to maximum localization.
+The size of the system is $256\times 256$ nodes.
+Now we move on to analyzing the stability of localization on single-node defect with disorder.
+For the topological QRW the disorder is added to rotation operators on the nodes in Eq.(2), via replacement $R(\Theta)\to R(\Theta+\delta\Theta)$ where $\delta\Theta\in[-\Theta_{dis},\Theta_{dis}]$ is a uniformly distributed random value centered around zero in finite interval.
+The defect node is not disordered.
+As an example of another mechanism, the localization of non-topological discrete-time QRW from Ref.
+[10] is studied for the same system sizes.
+The corresponding time step operator is defined as:
+$U=T_{y}C_{y}T_{x}C_{x},\quad T_{i}|\pm\rangle|i\rangle=|\pm\rangle|i\pm 1\rangle.$ (14)
+In this case the disorder is implemented as phases in the coin rotation operators:
+$C_{x}(\delta\Theta)=\frac{1}{\sqrt{2}}\begin{pmatrix}1&i\exp(i\delta\Theta)\\ i\exp(-i\delta\Theta)&1\end{pmatrix},$ (15)
+$C_{y}(\delta\Theta)=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-i\exp(i\delta\Theta)\\ -i\exp(-i\delta\Theta)&1\end{pmatrix}.$ (16)
+While different selections of disorder could be made, the one introduced above is motivated by the possible loss of phase coherence on the nodes of large system (described by slightly different phases acquired by passing each node).
+The sizes of the test systems are taken to be the same - $512\times 512$ nodes for single configuration of disorder, and $256\times 256$ for averaging over $20$ configurations.
+Both sizes are well above the saturation time threshold for example walker parameters ‘1’ and ‘2’ described in the previous sections.
+The results of simulation for different sizes of interval distributions $\delta\Theta\in[-\Theta_{dis},\Theta_{dis}]$ of angle disorder and for single configuration in coordinate space are presented in Fig.
+5.
+The values $\Theta_{dis}$ are equal for topological and non-topological QRWs.
+The disorder reduces the height of localization probability peak for both quantum walks, and consequently the corresponding efficiency of search algorithms, as it is shown in panels (a)-(d) for $\Theta_{dis}$ varying from $0.05$ to $0.5$..
+However, for one of the two topological walkers, the search probability always remains higher that the classical search probability during the first period of oscillation.
+Notably, the probability of quantum search (non-topological QRW) drops dramatically as the disorder increases.
+While the example shown in Fig.
+5 constitutes one specific disorder configuration and type, the trapped states demonstrate relatively stable behavior, even locally outperforming quantum search for large disorder.
+Still, we point out that the more systematic study should be made to define the optimal performance conditions of different quantum and classical search algorithms.
+One should note, however, that our algorithm still outperforms classical search in localization probability maximum at fixed time.
+The classical search always behaves as $P_{classical}(t)=t/N$ with $t$ number of check ups of individual nodes.
+Finally, the three panels of Fig.6 demonstrate a study of disorder effect on two mechanisms of localization with average over 20 spatial configurations.
+We show three different values of disorder interval limit $\Theta_{dis}\in\{0.3,0.5,1.0\}$.
+The small oscillations of the localization probability are not present in these plots, and the curves are more smooth.
+The effect of reducing the maximum value of $P_{localization}$ is present, especially for large disorder with $\Theta_{dis}=1$ in panel (c).
+The localization of non-topological QRW demonstrates a higher probability than both example walkers from topological QRW.
+V Conclusions
+In conclusion, we demonstrated that a fast dynamical localization phenomenon takes place in the split-step discrete-time topological quantum random walks.
+It was shown that localization mechanism is based on just a two pairs of localized trapped states associated with defect.
+Each pair of trapped states is particle-hole symmetry in the quasienergy spectrum.
+The time of highest localization is defined by the gap that separates two positive-energy trapped states.
+Due to a finite radius of localization, the position of these states in spectrum is not influenced in large enough systems.
+We should note that the overlap product criteria served as efficient tool to identify these trapped states, capture the evolution of walker and uncover the principle behind distribution of high-localization peaks in parameter space of walker.
+The described phenomena open an avenue towards realization of the search algorithms based on topological QRWs as well as further studies on explainability of localization phenomena in quantum random walks.
+The search mechanism, as formulated in this work, can be immediately tested in photonic and synthetic lattice quantum walk experiments that realize topological QRW [25, 26, 27, 28].
 Acknowledgements.
-We benefited from discussions with Anton Akhmerov, Yaroslav Herasymenko, Margarita Davydova and Carlo Beenakker.
+We benefited from discussions with Tereza Vakhtel, Tomas Osterholt, Anton Akhmerov, Yaroslav Herasymenko, Margarita Davydova and Carlo Beenakker.
 This project was supported by the Kavli Foundation.
@@
 This program (registration number 024.005.006) is part of the Gravitation program financed by the Dutch Ministry of Education, Culture and Science (OCW).
-Supplemental material for “Constant search time algorithm via topological quantum walks”
-by D. O. Oriekhov, Guliuxin Jin, and Eliska Greplova
-Appendix S1 Properties of split-step quantum random walk
+Code and data
+The code to reproduce research in this paper can be found at this URL: https://gitlab.com/QMAI/papers/topologicalsearch.
+The Zenodo repository containing both code and data can be found at this URL: https://zenodo.org/records/14853734.
+Appendix A Properties of split-step quantum random walk
 In this section we present a number of properties for split-step QRW that realizes topological phases in the case without defect nodes.
 Using Eq.(1) from the main text, and assuming the infinite system in both directions, we find the following equation for spectrum of translation-invariant unit cell:
-$\lambda^{2}+2\lambda\cos\left(\Theta_{2}/2\right)\left(\sin^{2}\left(\Theta_{1}/2\right)\cos\left(2k_{y}\right)-\cos^{2}\left(\Theta_{1}/2\right)\cos\left(k_{x}\right)\right)+\lambda\sin\left(\Theta_{1}\right)\sin\left(\Theta_{2}/2\right)\left(\cos\left(k_{x}-2k_{y}\right)+1\right)+1=0.$ (S1)
+$\lambda^{2}+2\lambda\cos\frac{\Theta_{2}}{2}\left(\sin^{2}\frac{\Theta_{1}}{2}\cos\left(2k_{y}\right)-\cos^{2}\frac{\Theta_{1}}{2}\cos\left(k_{x}\right)\right)$
+$+\lambda\sin\left(\Theta_{1}\right)\sin\frac{\Theta_{2}}{2}\left(\cos\left(k_{x}-2k_{y}\right)+1\right)+1=0.$ (17)
 The solution of this equation, $\lambda_{\pm}=e^{iE}$, gives a particle-hole symmetric spectrum of quasi-energies. The quasi-energy is given by
 $E=\pm\arccos\frac{b}{\sqrt{2b^{2}-4}},$
-$b=2\cos\left(\Theta_{2}/2\right)\left(\sin^{2}\left(\Theta_{1}/2\right)\cos\left(2k_{y}\right)-\cos^{2}\left(\Theta_{1}/2\right)\cos\left(k_{x}\right)\right)+\sin\left(\Theta_{1}\right)\sin\left(\Theta_{2}/2\right)\left(\cos\left(k_{x}-2k_{y}\right)+1\right)$ (S2)
+$b=2\cos\frac{\Theta_{2}}{2}\left(\sin^{2}\frac{\Theta_{1}}{2}\cos\left(2k_{y}\right)-\cos^{2}\frac{\Theta_{1}}{2}\cos\left(k_{x}\right)\right)$
+$+\sin\Theta_{1}\sin\frac{\Theta_{2}}{2}\left(\cos\left(k_{x}-2k_{y}\right)+1\right)$ (18)
 The topological phases and corresponding Chern numbers were defined in Ref.[11].
 The phase separation lines are characterized by the gap closing.
-For the technical purposed of identifying different types of search speed up discussed in the main text, we plot a map of gap value: the minimal gap value between two bands at $E=0$ or $E=\pm\pi$ (see Fig.S1(a) ) and the maximal gap value around the same energies.
+For the technical purposed of identifying different types of search speed up discussed in the main text, we plot a map of gap value: the minimal gap value between two bands at $E=0$ or $E=\pm\pi$ (see Fig.7(a) ) and the maximal gap value around the same energies.
 From such maps we find that the full gap closing happens only at lines $\Theta_{2}=2\pi n$ with $n$ being integer.
-$(a)$Refer to caption $(b)$Refer to caption
-Figure S1: The regions of one gap closing in panel (a) and both gaps closing in panel (b). Note the agreement between panel (a) and phase separation lines position.
-The colormap of search dependence on defect parameters for the walker ‘1’ shown in Fig.2 of the main text is presented in Fig.S2.
+Figure 7: The regions of one gap closing (upper panel) and both gaps closing (lower panel). Note the agreement between the upper panel and phase separation lines position in Fig. 2a.
+The colormap of search dependence on defect parameters for the walker ‘1’ shown in Fig.2 of the main text is presented in Fig.8.
 The dependence on $\Theta_{2}^{def}$ is absent, which motivates the studying of only different $\Theta_{1}^{def}$ values in the main text.
-Figure S2: @D colormap of dependence on defect parameters for a walker ‘1’. The optimal search is performed on a line - which demonstrates that the dependence on $\Theta_{2}^{def}$ is absent.
-Appendix S2 The comparison of product of overlaps with optimal search regions
-In this section we analyze how precise the correspondence between maximization of $|<d|n><n|i>|^{2}$ and $(\sum_{n=max_{1},max_{2}}|<d|n><n|i>|)^{2}$ is compared to the regions of optimal search in parameter space for a fixed defect parameters.
-For this purpose we compare the max search probability curves plotted along two lines depicted as dashed cyan in the first panel of Fig.S3.
+Figure 8: 2D color map of dependence on defect parameters for a walker ‘1’. The optimal search is performed on a line - which demonstrates that the dependence on $\Theta_{2}^{def}$ is absent.
+Appendix B The comparison of product of overlaps with optimal localization regions
+In this Appendix we analyze how precise the correspondence between maximization of $|<d|n><n|i>|^{2}$ and $(\sum_{n=max_{1},max_{2}}|<d|n><n|i>|)^{2}$ is compared to the regions of optimal localization in parameter space for a fixed defect parameters.
+For this purpose we compare the max search probability curves plotted along two lines depicted as dashed cyan in the first panel of Fig.9.
 One of these lines is a phase separation line and second one is parallel to phase separation and contains a walker `$1^{\prime}$ discussed in the main text.
 This second line crosses different topological phases.
-In the second and third panels of Fig.S3 we plot the comparison of max probability found on time interval $T=1000$ for the system size with $\sqrt{N}=40$ with values of different overlaps.
-The second and third panels in Fig.S3 clearly demonstrate that the combination of maximum and second maximum products of overlaps - the contribution of two pairs of trapped states - describes the probability peak most precisely.
+In the second and third panels of Fig.9 we plot the comparison of max probability found on time interval $T=1000$ for the system size with $\sqrt{N}=40$ with values of different overlaps.
+The second and third panels in Fig.9 clearly demonstrate that the combination of maximum and second maximum products of overlaps - the contribution of two pairs of trapped states - describes the probability peak most precisely.
 This leads to a conclusion that efficient search is performed by two pairs of trapped states in this system in all cases studied.
-Figure S3: Comparison of maximal product of overlaps described in the main text found over all eigenstates of evolution operator $\bar{U}$ for a two lines in parameter space plotted as dashed cyan lines in the first panel.
+Figure 9: Comparison of maximized over time localization probability with maximal product of overlaps described in the main text found over all eigenstates of evolution operator $\bar{U}$ for a two lines in parameter space plotted as dashed cyan lines in the first panel.
 The second panel corresponds to a line going through walker `$1^{\prime}$ and the third panel corresponds to a phase separation line.
 Next, we analyze the structure of trapped states and their position in spectrum for the example walker parameters labeled as `$2^{\prime}$.
-The analysis is similar to what is done in the main text and is presented in three panels of Fig.S4: comparison of search probability with overlap criteria in the first panel, the set of overlaps for each eigenstate of $\bar{U}$ in second panel and the comparison of probability evolution at the defect with the one defined by the two pairs of the trapped states in the third panel.
+The analysis is similar to what is done in the main text and is presented in three panels of Fig.10: comparison of search probability with overlap criteria in the first panel, the set of overlaps for each eigenstate of $\bar{U}$ in second panel and the comparison of probability evolution at the defect with the one defined by the two pairs of the trapped states in the third panel.
 However, in this case we find that the two pairs of trapped states that mainly contribute to the efficient search are placed near zero energy.
 Also there is no gap in spectrum around this energy, but the in-gap trapped states placed around $E=\pm\pi$ do not contribute into the constructive interference relevant for successful search.
-Figure S4: (a) The comparison of search probability with overlap criteria for walker ‘2’.
-(b) The identification of states that mainly contribute into constructive superposition on a searched node for walker ‘2’ parameters.
-(c) The comparison of density evolution at searched defect node (blue solid line) with the superposition (red dashed line) of trapped states identified from panel (b).
-Appendix S3 Structure of trapped states in coordinate space
+Figure 10: (a) The comparison of maximized over time localization probability with overlap criteria for walker ‘2’, with all quantities being normalized by their maximum value on the same grid of $\Theta_{1}^{def}$ parameters.
+(b) The identification of states that mainly contribute into constructive superposition on a searched defect node for walker ‘2’ parameters.
+(c) The comparison of density evolution at defect node (blue solid line) with the superposition (red dashed line) of trapped states identified from panel (b).
+Appendix C Structure of trapped states in coordinate space
 In this section we analyze how extended are the trapped states that mainly contribute to the search algorithm.
 From the search time dependence on system size shown in Fig.2(b) in the main text, we conclude that the more extended trapped state has radius $R\approx 50$ for walker ‘1’ and $R\approx 100$ for walker ‘2’ and defect parameter $\Theta_{1}^{def}=5\pi/8.$.
-The radius of trapped state was extimated as twice smaller that the system size where transition in time asymptotic happens.
-To further analyze the trapped state radius dependence on system size, we plot it in Fig.S5.
+The radius of trapped state was estimated as twice smaller that the system size where transition in time asymptotic happens.
+To further analyze the trapped state radius dependence on system size, we plot it in Fig.11.
 The set of tested system sizes covers the region where the trapped states are having overlap with itself due to periodic boundary condition.
@@
 Thus, the change in search time asymptotic is mostly defined by this second pair of trapped states, that has second maximal overlap value.
-Figure S5: Radius of trapped state as a function of system size.
-Left panel corresponds to a trapped state with largest overlap $|\langle d|n\rangle\langle n|i\rangle$ in Fig.3(b) of the main text, and right panel corresponds to the state with second largest overlap.
-Appendix S4 Stability of search algorithm with disorder
-In this section we make a stability comparison of the proposed algorithm of search based on trapped states with Grover’s algorithm in split-step quantum random walk described in Ref.[7].
-The size of the systems is taken to be the same - $512\times 512$ nodes, well above the saturation time threshold for both example walker parameters described in the main text.
-Since the Grover’s algorithm relies on the coherent superposition of most states present in the system, one could expect that disorder break such constructive superposition faster that for a few trapped states.
-To check that, we implement the angle disorder of spin rotation matrices on each node.
-For equation (2) in the main text such disorder is introduced as
-$R(\Theta)\to R(\Theta+\delta\Theta_{xy}),\quad\delta\Theta\in[-\delta\Theta_{max},\delta\Theta_{max}],$ (S3)
-where the distribution of $\delta\Theta_{xy}$ is uniform in the interval $[-\delta\Theta_{max},\delta\Theta_{max}]$.
-Similarly, for the split-step QRW analyzed in Ref.[7] and used here as Grover’s algorithm benchmark, the disorder is implemented as
-$C_{x}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&i\\ i&1\end{pmatrix}\to C_{x}(\delta\Theta)=\frac{1}{\sqrt{2}}\begin{pmatrix}1&i\exp(i\delta\Theta)\\ i\exp(-i\delta\Theta)&1\end{pmatrix},$ (S4)
-$C_{y}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-i\\ -i&1\end{pmatrix}\to C_{y}(\delta\Theta)=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-i\exp(i\delta\Theta)\\ -i\exp(-i\delta\Theta)&1\end{pmatrix},$ (S5)
-The results of simulation for equal intervals of disorder angle are presented in Fig.S6.
-The role of disorder in reducing efficiency of search algorithms in visible in all plots.
-However, for one of the two walkers, the search probability is always staying higher that the classical probability on the first period of oscillation.
-Notably, the probability of Grover’s search drops dramatically.
-In the presented example of disorder configuration and type, the trapped state are demonstrating relatively stable behavior as a basis of search algorithm, even outperforming Grover’s example for large disorder.
-Still, we point out that the more systematic study should be made to define the best performance conditions of different suboptimal quantum algorithms and compared with Grover’s one in noisy quantum setups.
-Also one should note that these algorithms still outperform classical search, which behaves as $P_{classical}(t)=t/N$ with $t$ being number of check ups of individual nodes.
-Figure S6: The search probability calculated for 2 walker parameters analyzed in the main text (see Fig.2) and compared with Grover’s algorithm in split-step QRW from Ref.[7] and with classical search with linearly increasing probability.
-The system size is $512\times 512$.
-Disorder strength varies from small in panel (a) $\delta\Theta_{max}=0.05$, to (b) $\delta\Theta_{max}=0.15$, (c) $\delta\Theta_{max}=0.35$, and large in panel (d) $\delta\Theta_{max}=0.5$.
-The walker $2$ shows best performance at large disorder out of all compared search procedures.
+Figure 11: Radius of trapped state as a function of system size.
+Left panel corresponds to a trapped state with largest overlap $|\langle d|n\rangle\langle n|i\rangle|$ in Fig.3(b) of the main text, and right panel corresponds to the state with second largest overlap.
+Two curves correspond to the radius of area in which $80\%$ and $95\%$ of state density concentrated.